OpenStudy (anonymous):

Find the slope and y intercept of the following 3x-2y=5

OpenStudy (aroub):

y=mx+b

OpenStudy (anonymous):

\[m=\frac{3}{2}\] \[y-intercept=\frac{5}{2}\]

OpenStudy (anonymous):

\[m=\frac{3}{2}\] \[y-intercept=-\frac{5}{2}\]

OpenStudy (anonymous):

my first answer was lacking a negative sign.

OpenStudy (aroub):

hy beehive can u please look at my question